OpenStudy (anonymous):

[6.05] Which set of equations would be used to solve this word problem? An airplane flew 3.5 hours with a 40 mph head wind. The return trip with a tail wind of the same speed took 2.5 hours. Find the speed of the plane in still air. d = 3.5(r + 40) and d = 2.5(r − 40) d = 3.5(r − 40) and d = 2.5(r + 40) d = 6(r + 40) None of these systems can be used to solve this problem.

OpenStudy (anonymous):

d = 3.5(r − 40) and d = 2.5(r + 40) => r=(3.5+2.5)*40=6*40=240